OpenStudy (anonymous):

A soccer team won 9 games and lost 3 games, what is the percentage of the total games played did the team win?

OpenStudy (lukecrayonz):

How many games total did they play?:)

OpenStudy (anonymous):

I'm assuming 12 total, 9 wins 3 losses. 75% of the total.

OpenStudy (lukecrayonz):

I know that.. :-) I was asking them

OpenStudy (lukecrayonz):

Find the total amount of games played.. 9+3=12 Take winning games, 9, and divide it by 12.. 9/12=3/4=75% :-)

OpenStudy (anonymous):

Yes exactly :)

OpenStudy (anonymous):

Thanks so much!